### v1

# A Simple Bias Reduction for Chatterjee’s Correlation

Abstract
Chatterjee’s rank correlation coefficient $\xi_{n}$ is an empirical index for detecting functional dependencies between two variables $X$ and $Y$. It is an estimator for a theoretical quantity $\xi$ that is zero for independence and one if $Y$ is a measurable function of $X$. Based on an equivalent characterization of sorted numbers, we derive an upper bound for $\xi_{n}$ and suggest a simple normalization aimed at reducing its bias for small sample size $n$. In Monte Carlo simulations of various cases, the normalization reduced the bias in all cases. The mean squared error was reduced, too, for values of $\xi$ greater than about 0.4. Moreover, we observed that confidence intervals for $\xi$ based on bootstrapping $\xi_{n}$ in the usual n-out-of-n way have a coverage probability close to zero. This is remedied by an m-out-of-n bootstrap without replacement in combination with our normalization method.
1 Introduction
The common correlation coefficients by Pearson, Spearman, or Kendall can only be used for identifying simple relationships between two random variables $X$ and $Y$. Pearson’s correlation coefficient approaches one only for a linear relationship, whereas Spearman’s $\rho$ and Kendall’s $\tau$ are based on ranks and thus less susceptible to the particular functional form of the relationship as long as it is monotone. For non-monotonous relationships, however, e.g. for $X$ uniformly distributed in $[-1,1]$ and $Y=X^{2}$, all three correlation coefficients can be zero on average.
Recently, Chatterjee has proposed a correlation coefficient $\xi_{n}$ that overcomes this restriction. He has shown that, as the data set size $n$ increases, this correlation coefficient converges to a theoretical quantity $\xi(X,Y)$ that is one, if $Y$ is a measurable function of $X$, and zero, if $X$ and $Y$ are independent [1]. The coefficient was subsequently generalized by Azadkia & Chatterjee [2] to a measure for conditional dependency and utilized in a model-free method for feature selection in machine learning (“FOCI”).
In case of independence, Chatterjee has shown that $\xi_{n}$ is asymptotically normally distributed with zero mean and standard deviation $2/5\sqrt{n}$ [1]. This can be utilized for an independence test, but, in a study by Shi et al. [3], this new test turned out to have less power than other independence tests. Hence, the main relevance of Chatterjee’s correlation coefficient is that it yields a normalized measure for the degree of dependence between two variables. Despite the asymptotic properties of the estimator $\xi_{n}$, we will prove that it is limited by an upper bound less than one and does not cover the possible range up to one. Therefore, $\xi_{n}$ can be considerably different from one, even in cases of strict linear relationship when other correlation coefficients are exactly one. For dependent $X$ and $Y$, this results in a systematic bias of Chatterjee’s estimator for $\xi$.
In the present article, we propose a simple rescaling of $\xi_{n}$ by dividing it through its upper bound in order to reduce its bias. In a series of Monte Carlo simulations, we measure the effect of this rescaling on the bias and mean squared error (MSE) of the estimator for $\xi$. Moreover, we consider the problem of constructing confidence intervals for $\xi$. As normality of $\xi_{n}$ only holds asymptotically and has only been proven by Chatterjee in the case of independence, we first computed confidence intervals based on the usual n-out-of-n bootstrap [4]. For continuous $Y$, however, we observed that the resulting intervals had a coverage probability close to zero and did not even include the point estimate. We therefore resorted to the m-out-of-n bootstrap, which has been shown to be more robust than the n-out-of-n bootstrap and to work in cases where the usual bootstrap fails [5].
Measuring bias, MSE, and coverage probability requires a computation of the true value of $\xi$. In order to be able to compute this value, we have transformed Chatterjee’s formula into an equivalent expression based on the conditional probability $P(Y\geq t|X=x)$, that further simplifies considerably for a continuous variable $Y$ (see Theorem 1 below).
2 Chatterjee’s correlation
Chatterjee’s correlation coefficient is based on the ranks of the values observed for $Y$. Let $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$ be $n$ pairs of sample values for $(X,Y)$, for which not all $y_{i}$ are identical, and which are ordered such that $x_{1}\leq x_{2}\leq\ldots\leq x_{n}$. Let $r_{i}$ be the rank of $y_{i}$, i.e., the number of values in $\vec{y}=(y_{1},\ldots,y_{n})$ that are less or equal to $y_{i}$. Additionally, let $l_{i}$ be the number of values in $\vec{y}$ that are greater or equal to $y_{i}$. Then Chatterjee’s rank correlation coefficient is defined as
$\xi_{n}(\vec{x},\vec{y})=1-\frac{n\sum_{i=1}^{n-1}|r_{i+1}-r_{i}|}{2\sum_{i=1}^{n}l_{i}(n-l_{i})}$ (1)
Note that in case of equal $x_{i}=x_{j}$ corresponding to different $y_{i}\neq y_{j}$, the value of $\xi_{n}$ depends on the tie breaking, which should be done at random according to Chatterjee. If all $x_{i}$ are different, this can not occur. Moreover, in the case of no ties among the $y_{i}$, the denominator is $n(n^{2}-1)/3$ and Eq. (1) simplifies to
$\xi_{n}(\vec{x},\vec{y})=1-\frac{3\sum_{i=1}^{n-1}|r_{i+1}-r_{i}|}{n^{2}-1}$ (2)
Chatterjee has shown that, if $(x_{i},y_{i})$ are i.i.d. pairs drawn according to the law of $(X,Y)$ and $Y$ is not almost surely constant, then $\xi_{n}(\vec{x},\vec{y})$ converges for $n\to\infty$ almost surely to the quantity
[…]
where the variable $t$ in the variance in the numerator is treated as a constant parameter, and $\mu$ is the probability distribution of $Y$. If $X$ and $Y$ are independent, the random variable $E(1_{\{Y\geq t\}}|X)$ is equal to $P(Y\geq t)$ and its variance is zero. If $Y$ is a measurable function of $X$, then $Y=f(X)$ is constant for a given event $X=x$, and $E(1_{\{Y\geq t\}}|X)=1_{\{Y\geq t\}}$, which means that numerator and denominator in (3) are identical.
In order to better understand expression (3) and make it easier to compute in practical use cases, we can rewrite it in terms of the conditional probability $P(Y\geq t|X=x)$ and the unconditional probability
[…]
where $\lambda$ is the probability distribution of $X$. First note that $E(1_{\{Y\geq t\}}|X)$ is just $P(Y\geq t|X)$. Further note that $1_{\{Y\geq t\}}$ is a Bernoulli variable which can take only two values:
$1_{\{Y\geq t\}}=\left\{\begin{array}[]{ll}1&\mbox{ with probability }P(Y\geq t)\\ 0&\mbox{ with probability }1-P(Y\geq t)\end{array}\right.$
The variance of a Bernoulli variable with $P(1)=p$ is $p(1-p)$, and, with the use of the formula $\mbox{Var}(Z)=E(Z^{2})-E(Z)^{2}$ for the variance in the numerator, Eq. (3) can be written as
$\xi$ $(X,Y)=$ (5)
[…]
where $\lambda$ is the probability distribution of $X$ and $\mu$ is the probability distribution of $Y$.
For a continuous random variable $Y$, two of the integrals can be solved symbolically, which results in a simplified formula for $\xi$ which was already proposed by Dette et al. as a dependency measure [6] [^1] Dette et al. used $\leq$” instead of $\geq$”, but, for continuous $Y$, this is equivalent as can be shown with the law of total probability..
Theorem 1.
For a continuous random variable $Y$,
[…]
Proof.
For continuous $Y$, we can write $d\mu(t)=f(t)dt$ with $f(t)=-\frac{d}{dt}P(Y\geq t)$. This yields
[…]
Inserting these values into (5) yields the theorem. Note that equalities (6) and (7) no longer hold for discrete random variables: In this case additional positive sums occur on the right hand side and the expressions are strictly greater than $1/3$, or $1/2$, respectively. ∎
The interesting property of $\xi$ with respect to dependency between $Y$ and $X$ can be readily seen in the expression in the above theorem. For independence of $X$ and $Y$, it is $P(Y\!\geq\!t|X=x)^{2}=P(Y\!\geq\!t)^{2}$ and the integral evaluates to $1/3$. If $Y$ is a function of $X$, i.e., $Y=f(X)$, $P(Y\!\geq\!t|X=x)^{2}$ is one if $f(x)\geq t$ and zero if $f(x)<t$, which means that the inner integral becomes $P(Y\geq t)$ and the outer integral evaluates to $1/2$.
3 Normalization of $\xi_{n}$
Although Chatterjee has proven in [1] that the coefficient $\xi_{n}$ converges almost surely to $\xi$, the average value of $\xi_{n}$ can be considerably different from the asymptotic value $\xi$ for small sample sizes. Consider, e.g., the extreme case of $X=Y$. In this case Pearson’s correlation is always one, whereas $\xi_{n}$ yields a considerably different value as the following sample R session with the use of the package XICOR [7] demonstrates:
> x <- runif(20)
> cor(x,x)
[1] 1
> xicor(x,x)
[1] 0.8571429
This is because the maximum possible value of Eq. (1) is less than one and this upper bound decreases as $n$ becomes smaller. To understand what the maximum value of $\xi_{n}(\vec{x},\vec{y})$ is over all possible permutations of the vector $\vec{x}$, we need the following
Lemma 1.
For real numbers $x_{1},\ldots,x_{n}$, it is
$\sum_{k=1}^{n-1}|x_{k+1}-x_{k}|\geq x_{max}-x_{min}$
where $x_{max}=\max\{x_{1},\ldots,x_{n}\}$ and $x_{min}=\min\{x_{1},\ldots,x_{n}\}$. Equality holds if and only if the $x_{i}$ are sorted in ascending or descending order.
Proof.
If $x_{min}=x_{max}$, the assertions are obvious. Let us assume that $i$ and $j$, $i\neq j$, are the indices with $x_{i}=x_{max}$ and $x_{j}=x_{min}$ and that, without loss of generality [^2] In the other case, we simply invert all indices $k\to n-k$., $i>j$. Then it is
$x_{max}-$ $x_{min}=|x_{i}-x_{j}|=|(x_{i}-x_{i-1})$
$\quad+(x_{i-1}-x_{i-2})\ldots+(x_{j+1}-x_{j})|$
$\leq\sum_{k=j}^{i-1}|x_{k+1}-x_{k}|\leq\sum_{k=1}^{n-1}|x_{k+1}-x_{k}|$
If the $x_{i}$ are sorted, equality holds, because for $x_{min}=x_{1}\leq x_{2}\leq\ldots\leq x_{n}=x_{max}$ it is11footnotemark: 1
$\sum_{k=1}^{n-1}|x_{k+1}-x_{k}|=\sum_{k=1}^{n-1}(x_{k+1}-x_{k})=x_{n}-x_{1}$
In order to prove that the reverse holds too, i.e., that if the sum takes its minimum, the $x_{i}$ are sorted, let us assume that the sum takes the minimum value $x_{max}-x_{min}$, that $x_{i}=x_{min}$ and $x_{j}=x_{max}$, and that [^3] In this case we will see that the sequence has ascending order. If $j>j$, we will have descending order and the proof remains the same, but with all indices inverted. $i<j$. Because of
$x_{max}-x_{min}$ $=\sum_{k=1}^{n-1}|x_{k+1}-x_{k}|$
$\geq\sum_{k=i}^{j-1}|x_{k+1}-x_{k}|\geq x_{max}-x_{min}$
equality holds and $x_{1}=x_{2}=\ldots=x_{i}$ and $x_{j}=x_{j+1}=\ldots=x_{n}$. Let us assume that the sequence is not sorted in ascending order, which means that there are indices $i^{\prime},j^{\prime}$ with $i\leq i^{\prime}<j^{\prime}\leq j$ and $x_{i^{\prime}}>x_{j^{\prime}}$. Applying the already proven inequality to different parts of the sum, we obtain
$x_{max}-$ $x_{min}=\sum_{k=i}^{j-1}|x_{k+1}-x_{k}|=\underbrace{\sum_{k=i}^{i^{\prime}-1}|x_{k+1}-x_{k}|}_{\geq\;x_{i^{\prime}}-x_{min}}$
$\quad+\underbrace{\sum_{k=i^{\prime}}^{j^{\prime}-1}|x_{k+1}-x_{k}|}_{\geq\;x_{i^{\prime}}-x_{j^{\prime}}}+\underbrace{\sum_{k=j^{\prime}}^{j-1}|x_{k+1}-x_{k}|}_{\geq\;x_{max}-x_{j^{\prime}}}$
$\geq x_{max}-x_{min}+2(x_{i^{\prime}}-x_{j^{\prime}})$
According to the assumption $x_{i^{\prime}}>x_{j^{\prime}}$, it is $02(x_{i^{\prime}}-x_{j^{\prime}})>$ 0, which is a contradiction, and the sequence must be in ascending order. ∎
With Lemma 1, it is easy to show that, for given values $\vec{x}$ and $\vec{y}$, $\xi_{n}$ takes its maximum value if the $r_{i}$ in Eq. (1) are sorted.
Theorem 2.
Let $\xi_{n}$ be defined as in Eq. (1). Then
$\xi_{n}(\vec{x},\vec{y})\leq\xi_{n}(\vec{y},\vec{y})$
where equality only holds for permutations of $\vec{x}$ that sort the $y_{i}$ in $(x_{i},y_{i})_{i=1}^{n}$ in ascending (or descending) order. The upper bound on the right is
$\xi_{n}(\vec{y},\vec{y})=1-\frac{n(n-r_{min})}{2\sum_{i=1}^{n}l_{i}(n-l_{i})}$
where $r_{min}$ is the absolute frequency of the smallest value in $\vec{y}$. When all $y_{i}$ are different, the upper bound is $(n-2)/(n+1)$.
Proof.
First note that the denominator in Eq. (1) only depends on the values $\vec{y}$ and is unaffected by a permutation of $\vec{x}$. Thus, $\xi_{n}$ is maximal if $\sum_{i}|r_{i+1}-r_{i}|$ is minimal. If the $r_{i}$ are in ascending or descending order, the value of this sum is $r_{max}-r_{min}=n-r_{min}$, which is the smallest possible value of the sum for all permutations of the $r_{i}$ according to Lemma 1. ∎
As the maximum value for $\xi_{n}$ according to Theorem 2 is less than one, it must be a biased estimator in case of $\xi(X,Y)=1$, i.e., a strict dependency between $X$ and $Y$. This can be remedied by rescaling $\xi_{n}$ to the full range up to one:
$\xi^{\prime}_{n}(\vec{x},\vec{y})=\frac{\xi_{n}(\vec{x},\vec{y})}{\xi_{n}(\vec{y},\vec{y})}$ (8)
As $\xi_{n}(\vec{y},\vec{y})$ approaches one for $n\to\infty$, $\xi^{\prime}_{n}$ is, like $\xi_{n}$, an asymptotically unbiased estimator for $\xi$. It is constructed in such a way that it has a smaller bias than $\xi_{n}$ if $\xi=1$, and in this case the mean squared error will be smaller, too. This does not necessarily hold, if the true value $\xi$ is less than one; in the special case of independent $X$ and $Y$, e.g., the MSE of $\xi^{\prime}_{n}$ will be greater because $\xi=0$ and $|\xi^{\prime}_{n}|\geq|\xi_{n}|$. The effect of the normalization (8) on the coverage probability of confidence intervals is yet another question that needs to be addressed. We therefore made extensive Monte Carlo simulations to measure these effects for different data generation processes.
4 Effect of normalization
For ten different models for the relationship between $X$ and $Y$ and for a range of data set sizes $n$, we have simulated $N=10^{6}$ data sets and compared the following quality indices for the estimators $\xi_{n}$ and $\xi^{\prime}_{n}$:
  • bias
  • mean squared error (MSE)
  • coverage probability of a $90\%$ confidence interval
The computation of $\xi_{n}$ has been done with the function xicor from the R package XICOR [7]. All three quality indices require knowledge of the true value $\xi$, which we have computed as described in the appendix. As the construction of confidence intervals for $\xi$ turned out to be more complicated than expected, we have devoted the separate Section 5 to this problem.
(a) model 1: $Y\sim X+\varepsilon$ (continuous)
(b) model 3: $Y\sim\sin(2\pi X)+\varepsilon$ (continuous)
(c) model 4: $Y\sim XY$ (discrete)
(d) model 6: $Y\sim X^{2}+\varepsilon$ (discrete)
Figure 1: Average value of $\xi_{n}$ and the normalized $\xi_{n}^{\prime}$ as a function of $n$ for different models. The solid black lines show the true value of $\xi$
4.1 Simulated models
The ten models include four models from Chatterjee’s original paper, namely three continuous models [1, Fig. 2] and the discrete model in [1, Sec. 4.2].
We use the notations unif$(a,b)$ for a uniform distribution between $a$ and $b$, norm$(\mu,\sigma^{2})$ for a normal distribution with mean $\mu$ and variance $\sigma^{2}$, equal$(a,b,n)$ for a discrete uniform distribution with $n$ values between $a$ and $b$, and binom$(n,p)$ for a binomial distribution with size $n$ and probability $p$.
Model 1:
$\boldsymbol{Y=X+\varepsilon}$
with $X\sim\mbox{unif}(-1,1)$ and $\varepsilon\sim\mbox{norm}(0,\sigma^{2})$.
This is the first model of Fig. 2 in [1].
Model 2:
$\boldsymbol{Y=X^{2}+\varepsilon}$
with $X\sim\mbox{unif}(-1,1)$ and $\varepsilon\sim\mbox{norm}(0,\sigma^{2})$.
This is the second model of Fig. 2 in [1].
Model 3:
$\boldsymbol{Y=\sin(2\pi X)+\varepsilon}$
with $X\sim\mbox{unif}(-1,1)$ and $\varepsilon\sim\mbox{norm}(0,\sigma^{2})$.
This is the third model of Fig. 2 in [1].
Model 4:
$\boldsymbol{Y=XZ}$
with $X\sim\mbox{binom}(1,p)$ and $Z\sim\mbox{binom}(1,p^{\prime})$.
This is the model used in [1] as an example for which the distribution looks normal even for dependent $X$ and $Y$. Chatterjee used the values $p=0.4$ and $p^{\prime}=0.5$.
Model 5:
$\boldsymbol{Y=X+\varepsilon}$
with $X\sim\mbox{equal}(m,-1,1)$ and $\varepsilon\sim-\sigma\sqrt{m^{\prime}}+\frac{2\sigma}{\sqrt{m^{\prime}}}\mbox{binom}(m^{\prime},0.5)$.
This is a discrete version of Model 1. $\sigma$ is the standard deviation of the noise $\varepsilon$. For $m$ and $m^{\prime}$, we have used the fixed values $m=6$, $m^{\prime}=2$.
Model 6:
$\boldsymbol{Y=X^{2}+\varepsilon}$
with $X\sim\mbox{equal}(m,-1,1)$ and $\varepsilon\sim-\sigma\sqrt{m^{\prime}}+\frac{2\sigma}{\sqrt{m^{\prime}}}\mbox{binom}(m^{\prime},0.5)$.
This is the discrete version of Model 2, where the noise is simulated by a multiple fair coin toss. The shift and the scaling factor are chosen in such a way, that $\varepsilon$ has zero mean and variance $\sigma^{2}$. For $m$ and $m^{\prime}$, we have used the fixed values $m=6$, $m^{\prime}=2$.
Model 7:
$\boldsymbol{Y=\sin(2\pi X)+\varepsilon}$
with $X\sim\mbox{equal}(m,-1,1)$ and $\varepsilon\sim-\sigma\sqrt{m^{\prime}}+\frac{2\sigma}{\sqrt{m^{\prime}}}\mbox{binom}(m^{\prime},0.5)$.
This is a discrete version of Model 1. $\sigma$ is the standard deviation of the noise $\varepsilon$. For $m$ and $m^{\prime}$, we have used the fixed values $m=6$, $m^{\prime}=2$.
Models 8-10:
$\boldsymbol{Y,X}$ independent
Here we examined one continuous and two discrete examples:
8)
$X\sim\mbox{unif}(-1,1)$ and $Y\sim\mbox{unif}(-1,1)$
9)
$X\sim\mbox{equal}(m,-1,1)$ and $Y\sim\mbox{equal}(m^{\prime},-1,1)$
10)
$X\sim\mbox{binom}(m,p)$ and $Y\sim\mbox{binom}(m^{\prime},p^{\prime})$
We have used $m=3,p=0.5$ and $m^{\prime}=6,p^{\prime}=0.3$
4.2 Effect on bias
For $\xi=0$, i.e., for independence of $X$ and $Y$, which we simulated in models 8-10, already $\xi_{n}$ itself showed to be unbiased. As $\xi^{\prime}_{n}$ is only a scaled version of $\xi_{n}$, it was unbiased, too, in these cases.
For $\xi>0$, however, the normalization reduced the bias in all cases. As the examples in Fig. 1 show, the normalization made the estimator almost unbiased even for small $n$ for models with a monotonous relationship between $X$ and $Y$, as in models 1 and 4. We observed the same for model 5 (not shown in Fig. 1). For more complicated relationships, as in models 2, 3, 6, and 7, the bias reduction was not as strong, but nevertheless quite distinct. This was to be expected because, even for strict functional relationship $Y=f(X)$, $\xi_{n}$ takes its maximum value only for monotonous $f$ (see Theorem 2). Normalizing by the maximum value thus can only partially correct for this bias.
(a) model 1: $Y\sim X+\varepsilon$ (continuous)
(b) model 2: $Y\sim X^{2}+\varepsilon$ (continuous)
(c) model 3: $Y\sim\sin(2\pi X)+\varepsilon$ (continuous)
(d) model 7: $Y\sim\sin(2\pi X)+\varepsilon$ (discrete)
Figure 2: MSE of $\xi_{n}$ and the normalized $\xi_{n}^{\prime}$ as a function of the true value of $\xi$ for different models and $n=30$.
4.3 Effect on MSE
As the normalization scales $\xi_{n}$ by a factor greater than one, it increases its variance. It thus happened that, for small $\xi$, the increase in variance outweighed the bias reduction for some models. Our models 1-3 and 5-7 allowed for covering the full range of $\xi$ by varying the parameter $\sigma$. The resulting MSE and its relation to the true value for $\xi$ for some fixed value of $n$ is shown in Fig. 2. As can be seen, the normalization typically reduced the MSE if $\xi\gtrsim 0.4$ and otherwise slightly increased it. The MSE reduction for high $\xi$ was greater, however, than its increase for small $\xi$.
5 Confidence intervals for $\xi$
The R function XICCOR::xicor() not only reports the value of $\xi_{n}$, but also a standard deviation based on Chatterjee’s asymptotic theory [1]. Unfortunately, the asymptotic theory has only been proven for independent $X$ and $Y$, i.e. for the special case $\xi=0$. Fig. 3 shows that the reported standard deviation becomes completely wrong as $\xi$ deviates from zero. It can therefore only be used to test the null hypothesis $H_{0}:\xi=0$, but not to construct parametric confidence intervals.
Figure 3: Comparison of the standard deviation of $\xi_{n}$ with the standard deviation reported by XICOR::xicor() for data simulated with model 1 and $n=30$.
(a) $X,Y\sim\operatorname{unif}(-1,1)$
(b) $X,Y\sim\operatorname{equal}(5,-1,1)$
Figure 4: Typical distribution (kernel density plot) of $\xi_{n}(\vec{x},\vec{y})$ for $n=50$, computed from n-out-of-n bootstrap samples $(\vec{x},\vec{y})^{*}$ for independent continuous (left) and discrete (right) $X$ and $Y$.
We therefore first tried the usual n-out-of-n bootstrap, which has been reported to provide non-parametric confidence intervals with decent coverage probability for a wide range of estimators [4, 8]. With $R=1000$ bootstrap repetitions, simulating a million cases was no longer feasible and we reduced the number of test cases to $N=10^{5}$. For all models with continuous $Y$, it turned out, however, that the bootstrap confidence intervals did not even include the point estimate $\xi_{n}$, which lead to coverage probabilities close to zero. This even occurred for independent $X$ and $Y$, as can be seen in Fig. 4(a). A possible explanation for this effect is that the bootstrap samples are not representative for the distribution of $Y$ because repeated drawing with replacement results in a discrete distribution with many multiple values, whereas in the original distribution of $Y$ a value almost never occurs multiple times. This means that the denominator in Eq. (1) is no longer $n(n^{2}-1)/3$ as it is for continuous $Y$. This explanation is in agreement with the absence of this effect for discretely distributed $Y$ (see Fig. 4(b)), although the coverage probability of the n-out-of-n bootstrap was unsatisfactory even in the discrete cases: For models 5-7 with $\sigma=0.1$ and sample size $n=2000$, the coverage probabilities of the usually best performing BCa bootstrap interval [4] were only about 0.65 even for the normalized estimator $\xi_{n}^{\prime}$, which was considerably less than the nominal value 0.9.
We therefore resorted to non-parametric intervals based on the m-out-of-n bootstrap, which has been shown to work in cases where the usual bootstrap fails [5]. For our problem, it is particularly appropriate because it allows for drawing without replacement and thus does not misrepresent a continuous distribution as a discrete distribution, as it happens by repeated drawing with replacement in the n-out-of-n bootstrap. Politis & Romano [9] have shown that the m-out-of-n bootstrap requires knowledge of an appropriate scaling factor $\tau_{n}$. With the aid of this factor, the quantiles $q(1-\alpha/2)$ and $q(\alpha/2)$ of the scaled bootstrap distribution $\tau_{m}(\xi_{m}^{*}-\xi_{n})$ are computed, where $\xi_{m}^{*}$ denotes the bootstrap samples obtained by m-fold drawing without replacement, and the confidence interval is estimated as
$\left[\xi_{n}-\frac{q(1-\alpha/2)}{\tau_{n}},\;\xi_{n}-\frac{q(\alpha/2)}{\tau_{n}}\right]$ (9)
Figure 5: Logarithmic plot of the variation of $\operatorname{Var}(\xi_{n})$ with $n$. The lines have been fitted with the model $\log\operatorname{Var}(\xi_{n})=\log V+\gamma\log n$.
In the subsequent article [10], Bertail, Politis & Romano explained that $\tau_{n}$ must be chosen such that $\tau_{n}^{2}\operatorname{Var}(\xi_{n})$ converges to some constant $V$. For the special case $\xi=0$, Chatterjee already proved this convergence for $\tau_{n}^{2}=n$ [1]. In the simulations of our models, we observed the same relationship $\operatorname{Var}(\xi_{n})\sim n^{-1}$, as can be seen in the examples in Fig. 5. Similar results were obtained for $\xi_{n}^{\prime}$. We therefore set the scaling factor to $\tau_{n}=\sqrt{n}$.
(a) model 1: $Y\sim X+\varepsilon$ (continuous)
(b) model 5: $Y\sim X+\varepsilon$ (discrete)
     model 3: $Y\sim\sin(2\pi X)+\varepsilon$ (continuous)
Figure 6: Coverage probability of 90% m-out-of-n bootstrap confidence intervals of the raw $\xi_{n}$ and the normalized $\xi_{n}^{\prime}$.
Another inconvenience of the m-out-of-n bootstrap is that it has a parameter $m$ which has to be chosen. For asymptotic convergence of the bootstrap distribution to the true distribution, Politis & Romano [9] proved the sufficient conditions $m\to\infty$ and $m/n\to 0$ for $n\to\infty$, but this still leaves a wide range of options, e.g. $m=c\cdot n^{\alpha}$ with $10<\alpha<$ 1. Moreover, they presented examples showing that the convergence rate of the coverage probability to the nominal value depends on the choice for $m(n)$.
# wrapper around xicor() for bootstrap
xiboot <- function(indices, data) {
  xicor(data[indices,"x"],
        data[indices,"y"]) /
        xicor(data[indices,"y"],
              data[indices,"y"])
# confidence interval after
# Politis & Romano (1994)
m.out.of.n.ci <- function(
     data, conf=0.90, R=1000)
  # point estimate
  xi.n <- xiboot(1:n, data)
  # bootstrap distribution
  m <- round(2*sqrt(n))
  indices <- replicate(R,
      sample(1:n, size=m, replace=F))
  xi.star <- apply(indices,
      MAR=2, xiboot, data=data)
  # confidence interval
  tau <- sqrt
  xq <- quantile(tau(m)*(xi.star-xi.n),
           c((1+conf)/2, (1-conf)/2))
  return(c(xi.n - xq[1] / tau(n),
           xi.n - xq[2] / tau(n)))
Listing 1: R implementation for computing the m-out-of-n bootstrap confidence interval for the normalized estimator $\xi_{n}^{\prime}$.
To circumvent this problem, different procedures for a data driven choice of $m$ have been suggested in the literature [11, 12, 13]. We have tried the method by Götze and Račkauskas [11], which minimizes the distance between the empirical cumulative distribution functions of the (scaled) statistic under consideration for the two m-out-of-n bootstraps with $m$ and $m/2$. As a distance measure, we have tried both the Kolmogorov distance and the $L_{2}$ distance, and we implemented a Golden Section Search [14] for efficiently finding the minimum. For all models, however, the resulting coverage probability was less than 0.8 over the range $50\leq n\leq 5000$.
We therefore tried different choices of the form $m=cn^{\alpha}$ with $c\in\{1,2,3,4\}$ and $\alpha\in\{1/2,2/3,3/4\}$. Among these combinations, the highest coverage probabilities for $500\leq n\leq 5000$ were achieved with $c=2$ and $\alpha=1/2$. We therefore made the choice $m=2\sqrt{n}$, and the code for computing the confidence intervals is given in Listing 1. As can be seen in Fig. 6, both the resulting coverage probability and the speed of the convergence to the nominal value was different for the different models. In all cases, however, it approached the nominal value for large $n$, and the coverage probability was higher for the normalized estimator $\xi_{n}^{\prime}$ than for the original estimator $\xi_{n}$. Moreover, for higher values of $\xi$, which correspond to smaller values for $\sigma$ in Fig. 6, the difference of the coverage probability was greater than for smaller values of $\xi$ (higher $\sigma$), and the confidence intervals for the normalized $\xi_{n}^{\prime}$ approached the nominal value 90% for much smaller $n$ compared to the use of the unnormalized $\xi_{n}$.
6 Conclusions
Our simulations show that scaling Chatterjee’s rank correlation coefficient $\xi_{n}$ by its upper bound both reduces its bias and improves the coverage probability of a non-parametric m-out-of-n bootstrap confidence interval. We have observed that the usual n-out-of-n bootstrap does not work, and our utilization of the m-out-of-n bootstrap without replacement, also known as “n choose m bootstrap”, is a viable approach to obtain asymptotically valid confidence intervals. For some of our simulated models, however, a nearly nominal coverage probability was achieved only for fairly high sample sizes $n\geq 1000$. It is thus an interesting subject for further research whether there are other ways to obtain a confidence interval for $\xi$ that approaches the nominal coverage probability already for smaller $n$ in more cases, or even in all cases.
Appendix: Computation of $\xi$
For estimating the bias of $\xi_{n}$ in the different models, it is necessary to compute the asymptotic value $\xi$. As $\xi_{n}$ is known to be asymptotically unbiased, the asymptotic value can be approximated by a Monte Carlo simulation with a large sample size. Alternatively, it can be done with Eq. (5) by means of symbolic integration, numeric integration, or a combination of both. In this appendix, the latter approach is chosen.
Model 1:
$\boldsymbol{Y=X+\varepsilon}$
with $X\sim\mbox{unif}(a,b)$ and $\varepsilon\sim\mbox{norm}(0,\sigma^{2})$.
In this case, the probability density of $X$ is $\mbox{dunif}(a,b)(x)=1/(b-a)$ for $x\in[a,b]$, and the probability density of $\varepsilon$ is $\mbox{dnorm}(0,\sigma^{2})(x)=\varphi(z/\sigma)/\sigma$, where $\varphi(z)=e^{-z^{2}/2}/\sqrt{2\pi}$ is the density of the standard normal distribution.
As $Y$ is the sum of two independent variables, its probability density $f_{Y}(y)$ is the convolution of the respective densities, i.e.,
[…]
$=\frac{1}{b-a}\left(\Phi\left(\frac{y-a}{\sigma}\right)-\Phi\left(\frac{y-b}{\sigma}\right)\right)$ (10)
where $\Phi$ is the cumulative distribution function of the standard normal distribution. Moreover,
[…]
$=\Phi\left(\frac{x-t}{\sigma}\right)$ (11)
The inner integral in Theorem 1 with respect to $d\lambda(x)=\mbox{dunif}(a,b)(x)\,dx$ can also be symbolically integrated to [15]
$\frac{1}{b-a}\int\limits_{a}^{b}P(Y\!\geq\!t|X\!=\!x)^{2}dx=\frac{\sigma}{b-a}\bigg{[}z\Phi(z)^{2}\\ +2\Phi(z)\varphi(z)-\frac{1}{\sqrt{\pi}}\Phi(z\sqrt{2})\bigg{]}_{\frac{a-t}{\sigma}}^{\frac{b-t}{\sigma}}$ (12)
Eventually, we have evaluated the remaining integral in Theorem 1 with respect to $d\mu(t)=f_{Y}(t)dt$ numerically with the R function integrate(). We used this mostly symbolic solution as a ground truth for testing the solely numeric integration that was necessary for the other models.
Models 2 & 3:
$\boldsymbol{Y=f(X)+\varepsilon}$
with $X\sim\mbox{unif}(a,b)$ and $\varepsilon\sim\mbox{norm}(0,\sigma^{2})$.
Both models can be written in this from with $f(x)=x^{2}$ for model 2, and $f(x)=\sin(x)$ for model 3. In this more general situation, it is
[…]
$=\Phi\left(\frac{f(x)-t}{\sigma}\right)$ (13)
and the other integrals cannot be symbolically solved. We therefore computed all other integrals with the function from Listing 2. It should be noted that the convolution integral py in Listing 2 might be computed more efficiently with an FFT based approach [16], but the computation by means of integrate() was fast enough for our purpose.
# numeric integration of xi
# for models f(X) + eps
#   px   = density function of x
#   peps = density function of eps
#   PYx  = P(Y>=t|x)
xi.cont.fx <- function(
        f, px, peps, PYx,
        minx=-Inf, maxx=Inf,
        miny=-Inf, maxy=Inf) {
  # convolution integral
  py <- function(y) {
    ff <- function(x) {
      px(x)*peps(y-f(x))
    integrate(ff, minx, maxx)\$value
  # integral w.r.t. dlambda(x)
  inner.integral <- function(t) {
    ff <- function(x) {
      PYx(t,x)^2 * px(x)
    integrate(ff, minx, maxx)\$value
  # integral w.r.t dmu(t)
  ff <- Vectorize( function(t) {
    py(t) * inner.integral(t)
  6*integrate(ff, miny, maxy)\$value - 2
Listing 2: R implementation of the numeric integration to compute the value $\xi$ according to Theorem 1.
Model 4:
$\boldsymbol{Y=XZ}$
with $X\sim\mbox{bernoulli}(p)$ and $Z\sim\mbox{bernoulli}(p^{\prime})$.
This is the model for which Chatterjee already gave a formula for $\xi$ [1]. $Y$ can only take the two values 00 and $1$ with probabilities
$P(Y=1)=pp^{\prime}\quad\mbox{and}\quad P(Y=0)=1-pp^{\prime}$ (14)
and we have
$P(Y\geq 0|X\!=\!x)$ $=1$ for both $\quad\mbox{for both }x$ (15)
$P(Y\geq 1|X\!=\!x)$ $=p^{\prime}$ for $\quad\mbox{for }x=1$
$P(Y\geq 1|X\!=\!x)$ $=0$ for $\quad\mbox{for }x=0$
Inserting everything into Eq. (5) yields
$\xi=\frac{(1-p)p^{\prime}}{1-pp^{\prime}}$ (16)
Model 5 & 6:
$\boldsymbol{Y=f(X)+\varepsilon}$
with $X\sim\mbox{equal}(a,b,n)$ and
$\varepsilon\sim-\sigma\sqrt{m}+\frac{2\sigma}{\sqrt{m}}\mbox{binom}(m,0.5)$.
For discrete distributions, all integrals in Eq. (5) actually are finite sums which can be readily evaluated by the computer. For computing the distribution of $Y$, i.e. the convolution between the distributions of $f(X)$ and $\varepsilon$, we have used the function conv() from the R package kSamples [17].